Jay Levitt Company budgeted the following cost standards for the current year:
Direct materials (2 kg of plastic at $6 per kilogram) $12.00
Direct labour (2 hours at $12 per hour) 24.00
Variable manufacturing overhead 11.90
Fixed manufacturing overhead 6.25
Total standard cost per unit $54.15

Actual costs for producing 2,740 units were as follows:
Direct materials used 5,560 kg
Direct materials purchased (6,780 kg) $40,002
Direct labour (6,880 hours) $67,424
Variable manufacturing costs $32,600
Fixed manufacturing costs $17,600
Partially correct answer. Your answer is partially correct. Try again.
Calculate the materials price variance for materials purchased.
Materials price variance $enter the materials price variance

Entry field with incorrect answer

select a type of the materials price variance

Entry field with correct answer UnfavourableFavourableNeither favourable nor unfavourable

Partially correct answer. Your answer is partially correct. Try again.
Calculate the materials quantity variance.
Materials quantity variance $enter the materials quantity variance

Entry field with incorrect answer

select a type of the materials quantity variance

Entry field with correct answer Neither favourable nor unfavourableFavourableUnfavourable

Partially correct answer. Your answer is partially correct. Try again.
Calculate the labour price variance. (Round answer to the nearest whole dollar, e.g. 5,275.)
Labour price variance $enter the labour price variance rounded to the nearest whole

Entry field with incorrect answer

select a type of the labour price variance

Entry field with correct answer UnfavourableFavourableNeither favourable nor unfavourable

Partially correct answer. Your answer is partially correct. Try again.
Calculate the labour quantity variance.
Labour quantity variance $enter the labour quantity variance

Entry field with incorrect answer

select a type of the labour quantity variance

Entry field with correct answer UnfavourableNeither favourable nor unfavourableFavourable
Answer #1

1) Material price variance = (Standard price-actual price)actual qty = (6*6780-40002) = 678 F

Material quantity variance = (2740*2-5560)*6 = 480 U

2) Labor rate variance = (12*6880-67424) = 15136 F

Labor efficiency variance = (2740*2-6880)*12 = 2800 u